A meteor is moving at a speed of 100ms^-1 towards mars is detected at a height of 20.7 m from its surface . The meteor reaches the surface after 2 seconds . What is the gravitational acceleration of mars?


1
Expert's answer
2021-01-27T09:30:30-0500

If the velocity vector is directed to the surface of Mars (to its center), then the condition is incorrect.

Assume that the velocity vector be directed at an angle of "86\u00b0" to the vertical. We get


"s=v_0\cdot \cos86°\cdot t+\frac{g_Mt^2}{2}\to"


"20.7=100\cdot2 \cdot\cos86°+g_M\cdot2^2/2\to g_M\approx3.4(m/s^2)" . Answer